OpenStudy (anonymous):

can someone plz tell me how to find the coordinates of center of mass using double integration(iterated integrals)?also plz give an example(for instance semi circular disc)?

OpenStudy (anonymous):

someone plz help!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

i think the formola goes like this \[\int\limits_{}^{}\int\limits_{}^{}\rho dm \div \int\limits_{}^{}\int\limits_{}^{}dm\]

OpenStudy (anonymous):

ok thanks!!! but what is x here

OpenStudy (anonymous):

x is the coordinate of the element dA. similarly integral y pho dA

OpenStudy (anonymous):

but how han i use it to find com of a semicircular disk

OpenStudy (anonymous):

http://cnx.org/content/m14120/latest/

OpenStudy (anonymous):

ok thanks very much but why isnt d double integral used here?

OpenStudy (anonymous):

coz by summetry x for com is zero. so, we need to do only for y

OpenStudy (anonymous):

ok so if we consider the case of a disc we have to use it r8?

OpenStudy (anonymous):

yes